Status Report

NASA Space Station On-Orbit Status 14 Mar 2004

By SpaceRef Editor
March 14, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 14 Mar 2004

All ISS systems continue to function nominally, except as noted previously or below.  Day 148 in space for Expedition 8 (146 days aboard ISS), and a Sunday.  Ahead: Week 21 of Increment 8.

The FE performed another daily session of inspecting the BIO-5 Rasteniya-2 (“Plants-2”) experiment which studies growth and development of plants (peas) under spaceflight conditions in the Lada-4 greenhouse.

The FE also performed the daily routine environmental control & life support systems (SOZh) maintenance in the Service Module (SM).

Mike and Sasha conducted their regular daily physical exercise program of 2.5 hrs on TVIS treadmill, CEVIS bike, RED exerciser and VELO cycle with load trainer.

Early this morning, CMG-2 (control moment gyroscope #2) suffered a failure.  The automated Failure Detection, Isolation, and Recovery system (FDIR) restored the gyro to operation during an LOS (loss-of-signal) period.   [The failure occurred at 00:12am EST, and recovery was accomplished two minutes later (00:14am).  At AOS (acquisition-of-signal), CMG performance was back to normal.  Zone-of-exclusion (ZOE) recorders captured technical performance data of the incident for later analysis.]

Weekly Science Update (Expedition Eight — 17th):

GASMAP:   Next GASMAP 30 Day Health Check activity is scheduled for 3/26.

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound (ADUM):   Mike Foale was thanked for completing the ADUM OPE (On-board Proficiency Enhancer) on Friday (3/12) and downlinking the text file.  The ground team hopes that he enjoyed the CD (compact disk) and feels confident about his ADUM scan session next week.

Hand Posture Analyzer (HPA:  Looking forward to future operations.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   Looking forward to future operations, after the PromISS operations.
Space Acceleration Measurement System (SAMS):   The Science Officer was thanked for rebooting the SAMS ICU (interface control unit) computer in support of capturing the TVIS IFM (inflight maintenance) acceleration data.
Microgravity Acceleration Measurement System (MAMS):   MAMS operations are nominal.

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Behaving nominally.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):   Experiment has ended.  The crew was thanked for all of their hard work and support with PromISS stowage activities.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Renal Stone (RS):   The crew’s last in-flight data collection session will occur in April.  Thanks to the crew for their continued support with taking their daily pills at dinner time.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   Pre-Increment requirements have been completed.  A second BBT (Beacon & Beacon Tester) session will be scheduled in US Lab.

Foot/Ground Reaction Forces during Space Flight (FOOT):   The Science Officer was thanked for his willingness to press through this timeline-intensive session and for his efforts in obtaining a maximum data take.  The ground team will provide results of this session as soon as the data’s on the ground.  Currently the data downlink is scheduled for next week, scheduled for the coming week.

Materials ISS Experiment (MISSE):  In progress. Deployed outside. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   Looking forward to the next set of FDI Tissue Culture runs.

Yeast Group Activation Packs (Yeast GAP):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  Nothing new. 

Earth Science Toward Exploration Research (ESTER):  Looking forward to the upcoming sessions.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   CGBA continues to run nominally at 20C.

Miscible Fluids in Microgravity (MFMG):    Nothing new.

Educational Payload Operations (EPO):  Education looks forward to the upcoming Flight Demonstration, which supports national science education standards.  The video will be used in museum and science center student and educator programs.

Crew Earth Observations (CEO):  An interesting ISS/CEO image of Iceberg A-39D of South Georgia Island has been posted on Earth Observatory this weekend (see link, below).  The National Snow and Ice Data Center believes that the unusual blue color of the western tip of the iceberg was due to varying depths of snowmelt pooled there during the Southern Hemisphere summer.  Also to be posted in the Dynamic Events section is one of your nice oblique views of last Monday’s Dust Storm Cloud band that moved off western Africa over the Atlantic as it was being swept northeastward in advance of an approaching cold front.  Thanks to the crew for their fine response to the ground’s request for imagery of this event.  Preliminary reviews of their most recent imagery indicate that they have acquired a very useful set of images of the investigators’ coral reef targets in the South Pacific.  They will be cataloging them soon.  Camera times and focus are looking good.

Today’s CEO (Crew Earth Observations) targets, limited in XPOP attitude by flight rule constraints on the use of the science window, which is available for only ~1/4 of each orbit when not facing forward (in “ram”), were Sydney, Australia (panorama looking left of the entire Sydney basin), Lake Eyre, Australia (long-term monitoring site.  Water levels in this low point in Australia’s vast interior non-outlet basin fluctuate on a multi-year periodicity that is poorly understood.  ISS passed over the main inflow point: looking at nadir and rightwards for the lake floor), and Chao lava dome, Chile (400mm-lens.  Chao is the largest viscous-lava dome in the world, 14.5 km long with steep lava walls 350-400 m high. It is less than 100,000 years old and has an erupted volume of ~26 cubic-km.  The ‘elephant skin’ texture on the surface is due to flow ridges, which are about 30m high.  The dome lies at 5000 m in the high western Andes Mts.  Nadir pass for requested detailed images).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

SpaceRef staff editor.